OpenStudy (anonymous):

determine the domain f(x)=sq root sign over 7-x

OpenStudy (anonymous):

\[f(x)=\sqrt{7-x}\]

OpenStudy (anonymous):

\[f(x)=\sqrt{7-x}\] you cannot take the square root of a negative number so solve \[7-x\geq 0\] get \[7\geq x\] is your domain

OpenStudy (anonymous):

as an interval it is \[(-\infty,7]\]